OpenStudy (anonymous):

A bacteria culture grows with constant relative growth rate. After 2 hours there are 400 bacteria and after 8 hours the count is 50,000. Find the rate of growth after 7 hours. P'(7) = ?

OpenStudy (anonymous):

equation follows as P(7) = 80e^(r(7))

OpenStudy (anonymous):

i need to take the derivative of P(7)

OpenStudy (gorv):

can u give what is initial equation of P

OpenStudy (anonymous):

P(0) = 80

OpenStudy (gorv):

thatz it ??

OpenStudy (gorv):

how did u get e^(r(7))

OpenStudy (anonymous):

from the growth rate formula of C*exp^(rt)

OpenStudy (gorv):

rt = rate ??

OpenStudy (gorv):

p(0)=80

OpenStudy (anonymous):

r = rate t = time

OpenStudy (gorv):

P(t)=80*e^(r*t)

OpenStudy (anonymous):

yes

OpenStudy (gorv):

we need to find r first

OpenStudy (gorv):

after 2 hour 400=80*e^2r ........(1)

OpenStudy (anonymous):

.8047

OpenStudy (gorv):

r=0.8047 u got ??

OpenStudy (anonymous):

yes

OpenStudy (gorv):

so we need to find P'(7) first find P(t)

OpenStudy (gorv):

oh i mean P ' (t)

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

natural log?

OpenStudy (gorv):

no no just differentiate

OpenStudy (gorv):

P(t)=80*e^(0.8047*t)

OpenStudy (gorv):

\[P'(t)=\frac{ d P(t) }{ dt }\]

OpenStudy (gorv):

P'(t)=0.8047*80*e^(0.8047*t)

OpenStudy (gorv):

now plug in t=7

OpenStudy (anonymous):

what rule did u use to differentiate?

OpenStudy (gorv):

\[P'(7)=0.8047*80*e^{0.8047*7}\]

OpenStudy (gorv):

u know the differentiation of e^2x ??

OpenStudy (anonymous):

it doesn't change right?

OpenStudy (gorv):

yeah but 2 will come froward

OpenStudy (gorv):

=2*e^2x

OpenStudy (gorv):

right ??

OpenStudy (anonymous):

i believe so haha, ill try out the answer i get when plugging in t and see if its right

OpenStudy (gorv):

similarly 0.8047 sill come forward

OpenStudy (anonymous):

it worked! thank you man

OpenStudy (gorv):

welcome :)

OpenStudy (paxpolaris):

\[P'(7)={125\over2}\cdot \ln \left( 5 \right)\cdot \sqrt5\]

OpenStudy (gorv):

how???

OpenStudy (gorv):

80*0.8047=64.3 125/2=62.5

OpenStudy (paxpolaris):

sorry ... 80 times that

OpenStudy (paxpolaris):

\[P'(7)={5000}\cdot \ln \left( 5 \right)\cdot \sqrt5 \]